Chad the bodybuilder and his buddies run a moving company. The company employs two types of workers: skinny guys (x₁) and beefy guys (x₂). The production function of the company is f(x₁,x₂)=x₁+2x₂, where x₁ is the number of skinny guys and x₂ is the number of beefy guys. If each skinny guy gets paid $1 and each beefy guy gets paid $3, what is the minimal cost of production of 50 units of output?
Answer #1

MRTS = MP1/MP2 = 1/2

MRTS is constant, so perfect Substitutes preferences,

so only corner solution exists

P1/P2 = 1/3

as MRTS > P1/P2

MP1/P1 > MP2/P2

so , only input 1, X1 is employed

X1* = f = 50

so total cost = P1*X1

= 50

So Minimum cost = $ 50